OpenStudy (anonymous):

what is the length of the hypotenuse of a right triangle that has legs with lengths of 5 and 4

OpenStudy (ganpat):

\[(hyp)^{{2}} = (side1)^{2} + (side2)^{2}...\]

OpenStudy (anonymous):

can u do it for me

OpenStudy (anonymous):

all my answers choices are in square root

OpenStudy (anonymous):

square root of 51 and 41 and 49 and 47

OpenStudy (ganpat):

\[5^{2} + 4^{2} = 25 + 16 = 41... so \sqrt{41}...\]

OpenStudy (anonymous):

yayyy thank you

OpenStudy (ganpat):

W/c